# Random walks on edge colored random graphs

### v2

@@ 1 Introduction
The cover time of a connected graph is the maximum over the start vertex of the expected time for a simple random walk to visit every vertex of the graph. There is a large literature on this subject see for example [1], [13], including [4][8] which give precise estimates of cover time for various models of random graphs.
@@
(Smooth walks) Let $G$ be the union of a random $r$-regular graph with red edges and a $b$-regular graph with blue edges where $r\geq 1$ and $b\geq 2$.
@@
    If $b\geq 3$ and $\alpha<1$ constant, then w.h.p. the cover time of an $\alpha$-constrained smooth random walk on $G$ with budget $\gamma(\alpha)n\log n$ where $\gamma(\alpha)=(\alpha\sigma_{RB}r)(r+b)$ is asymptotically equal to $C_{G}(R,B)=(\alpha\sigma_{RB}+(1-\alpha)\sigma_{B})n\log n$.
@@
In a flip walk there is no hard bound on the number of times we use a red edge, only a stochastic one. At any step, let $\rho_{R}$ be the probability of transition down a given red edge adjacent to the current vertex and $\rho_{B}$ the probability of transition down a given blue edge. Thus the probability the next transition is red is $r\rho_{R}$ etc. Examples include the following. Let $\rho_{R}=\alpha/r$, and $\rho_{B}=(1-\alpha)/b$ in which case the probability of transition down a red edge at any step is $\alpha$, and a blue edge is $1-\alpha$.
@@
    Suppose $b\geq 3$. and the budget is $(1-\varepsilon)\sigma_{RB}n\log n$.The cover time of an oblivious walk on $G$ is asymptotically equal to $(\alpha\sigma_{RB}+(1-\alpha)\sigma_{B})n\log n$ w.h.p., as given in (1) where $\alpha=\tfrac{1-\varepsilon}{r(r+b)}$.
@@
    Suppose that $b=2$, $r\geq 1$, the blue edges span a Hamilton cycle, and the budget be at most $(1-\varepsilon)(r+1)/rn\log n$. The cover time of an oblivious walk on $G$ is asymptotic to $C_{G}(R,B)\sim n^{2}/2$, w.h.p.
@@
The main tool in our proofs will be the First Time Visit Lemma, a technique which will allow us to make very precise estimates of cover time in certain classes of graphs. The technique was introduced in [4], and subsequently refined in [7], and in [8] where various technical conditions were removed. In outline, the approach is to bound the cover time above and below using (4) of Lemma 4. Much of the work is in constructing the lower bound on the cover time. On regular graphs the walks we consider have uniform stationary distribution, and are rapidly mixing. The unvisit probability (4) has a parameter $p_{v}\sim\pi_{v}/R_{v}$ (see (3)) where $R_{v}$ is the expected number of returns to vertex $v$ during the mixing time.
@@ 2 Background material
We use $A_{n}\sim B_{n}$ to denote $A_{n}=(1+o(1))B_{n}$ and thus $\lim_{n\rightarrow\infty}A_{n}/B_{n}=1$. We use $\omega$ to denote a quantity which tends to infinity with $n$ more slowly than any other variables in the given expression. The expression $f(n)\ll g(n)$ indicates $f(n)=o(g(n))$. The inequality $A\lesssim B$ which stands for $A\leq(1+o(1))B$ is used to unclutter notation in some places.
@@ 2.1 First Time Visit Lemma
We denote the steady state probability of a simple random walk on $G=(V,E)$ being at $v\in V$ is $\pi_{G}(v)$. We let $\mathbb{P}_{u,G}^{(t)}(v)$ denote the probability that a simple random walk on $G$ started at $u$ is at $v$ after $t$ steps. We then define a mixing time $T_{G}$ by
@@
For $t\geq 0$, let ${\mathcal{A}}_{t}(v)$ be the event that ${\mathcal{W}}_{u}$ does not visit $v$ in steps $T_{G},T_{G}+1,\ldots,t$. The vertex $u$ will have to be implicit in this definition. We will use the following version:
@@
Let $R_{v}\geq 1$ is the expected number of visits by ${\mathcal{W}}_{v}$ to $v$ in the time interval $[0,T_{G}]$. Suppose that $T_{G}\pi_{v}=o(1)$ and $R_{v}=O(1)$. Let
@@ 3.1 Proof of Theorem 1(a)
In this section we prove that if $b\geq 3$ then w.h.p. the cover time of an $\alpha$-constrained smooth random walk on $G$ with budget $\gamma(\alpha)=(\alpha\sigma_{RB}r)/(r+b)\,n\log n$ is asymptotically equal to $C_{G}=(\alpha\sigma_{RB}+(1-\alpha)\sigma_{B})\,n\log n$. It was shown in [4] that the unconstrained cover time of $G$ is $\widehat{C}_{G}\sim\sigma_{RB}\,n\log n$ w.h.p. Without any constraint on the choice of the next edge, a walk of length $\widehat{C}_{G}$ would cross a red edge $\sim\frac{r}{r+b}\sigma_{RB}\,n\log n$ times. It follows that w.h.p. we will not be able to complete the walk within the budget without having to avoid red edges a significant number of times.
If the phases are $o(T_{G})$ in length then there are some technical difficulties that we have not yet managed to overcome, see Remark 1 below.
@@
There does not seem to be a compelling reason for us to only consider smooth walks. If we allow the walk to switch between allowing red edges and not allowing them, then we have no problems with mixing. This is because the steady state probabilities are $1/n$ regardless. The problem arises because $R_{v}$ will depend heavily on how we switch between the two possibilities.
@@ 3.1.1 Upper bound
Suppose now that we start a walk at vertex $u$ and that we alternate between red/blue phases and blue phases. Take $t_{0}=0$, and suppose we change phase at times $t_{1},t_{2},\ldots,t_{i},\ldots$. Assume that whenever $i$ is even we start a red/blue phase at $t_{i}$ and switch to a blue phase when $i$ is odd. At each step in a red/blue phase there is an $r/(r+b)$ chance of using a red edge. We also know that this walk will finish at least as early as one that only uses blue edges. This follows from Lemma 4 and the fact that $s/(s-1)$ decreases as $s$ increases. We will assume $t_{i}-t_{i-1}\geq\omega\log^{2}n\gg T_{G}$ for $1\leq i<m$, so that
$t_{RB}=\sum_{i\geq 0}(t_{2i+1}-t_{2i})\leq\left(1-\Theta\left(\frac{\log n}{n^{1/2}}\right)\right)\alpha\sigma_{RB}n\log n.$ (6)
Multiplying the RHS of (6) by $r/(r+b)$ and applying the Chernoff bounds yields an upper bound of $\gamma(\alpha)n\log n$ on the number of times a red edge is used.
@@ 3.1.2 Lower bound
For the lower bound we let $t_{1}=\tfrac{(1-\varepsilon)\alpha(r+b)}{(b-2)(r+b-1)}n\log n$ and argue that $C_{u}\geq t_{1}$ w.h.p. Proceeding as in [4], for any vertex $u$, we can find a set of vertices $S$ such that at time $t_{1}$, the probability the set $S$ is covered by the walk ${\mathcal{W}}_{u}$ tends to zero. Hence $T_{G}(u)>t_{1}$ whp which together with (13) completes the proof of Theorem 1(a).
We construct $S$ as follows. Let $\sigma$ be given by (5). Let $S\subseteq[n]$ be some maximal set of locally tree-like vertices all of which are at least distance $12\sigma+$ 1 apart. Thus $|S|\geq(n-r^{3\sigma})/r^{(2\sigma+1)}$.
@@
$\mathbb{P}(S(t_{1})\neq 0)\geq\frac{(\mathbb{E}|S(t_{1})|)^{2}}{\mathbb{E}|S(t_{1})|^{2}}=\frac{1}{\frac{{\bf E}|S_{t_{1}}|(|S_{t_{1}}|-1)}{({\bf E}|S(t_{1})|)^{2}}+(\mathbb{E}|S_{t_{1}}|)^{-1}}=1-o(1).$
@@
There is a natural measure preserving mapping from the set of walks in $G$ which start at $u$ and do not visit $v$ or $w$, to the corresponding set of walks in $\Gamma$ which do not visit $Z$. Thus the probability that ${\mathcal{W}}_{u}$ does not visit $v$ or $w$ in the first $t$ steps is equal to the probability that a random walk $\widehat{{\mathcal{W}}}_{u}$ in $\Gamma$ which also starts at $u$ does not visit $Z$ in the first $t$ steps.
@@ 4 Proof of Theorem 2
where $E(S:S^{c})$ is the number of edges between $S$ and $S^{c}$ and ${\Phi}_{SRW}$ refers to the conductance of a simple randpm walk.
As $G$ satisfies P5, the conductance ${\Phi}^{*}_{SRW}$ of a simple random walk on $G$ is constant for some $10<c<$ 1. Assuming $\min(\rho_{R},\rho_{B})$ is also a positive constant, the conductance ${\Phi}^{*}_{PB}$ of the flip walk is constant, i.e.,
@@
Consider an infinite $d$-regular tree $T_{d}$, where $r+b=d$ and in which each vertex has $r$ red and $b$ blue edges. For discussion we regard $T_{d}$ as rooted at a designated vertex $v$, with all edges nominally oriented outward from $v$. With this orientation, a transition away from $v$ (resp. $u$) at a vertex $u$, is a transition $uw$ down an out-edge of $u$. Relative to this orientation, a vertex other than $v$ is red if its in-edge is red, and blue if its in-edge is blue. With the exception of $v$ this induces an alternating pattern of red and blue vertices in $T_{d}$, in which e.g., a red vertex has $(r-1)$ red out-edges, and $b$ blue out-edges.
@@
$\phi_{RR}=$ ( $\mathbb{P}(\text{ the walk returns to }u\mid u\text{ is red and the walk exits away from }u\text{ on a red edge}),$
and define $\phi_{RB},\phi_{BB},\phi_{BR}$ similarly. Note that $\phi_{RR}=\phi_{BR}=\phi_{v,R}=\phi_{R}$, say. For a red vertex $u$, let $\psi_{R}(u)$ be the probability the walk moves away from $v$ at $u$, and subsequently returns to $u$. Thus with $P(u,w)$ the colour dependent transition probability over the edge $uw$, $\phi(uw)$ the colour dependent return probability, and $N^{+}_{R}(u)$ the red out-neighbours of $u$ etc.,
@@
If $u$ is red, there are $r-1$ red out-neighbours, and $b$ blue out-neighbours, so
@@
Put $\psi_{R}=x,\psi_{B}=z$. As $r=1$ we have $\rho_{R}+b\rho_{B}=1$. Assume $b\geq 2$ and $\rho_{R}\leq\rho_{B}$ so that $0\leq\rho_{R}\leq 1/(b+1)$. Use (19) to get
$\frac{1}{1-x}=\frac{1-z}{1-z-b\rho_{B}^{2}},$
and substitute this for $1/(1-x)$ in (20) giving
$z=\frac{\rho_{R}^{2}(1-z)}{1-z-b\rho_{B}^{2}}+\frac{(b-1)\rho_{B}^{2}}{1-z}.$
Put $q=b\rho_{B}$ and $\rho_{R}=1-q$ where $b/(b+1)\leq q\leq 1$, to consider $F(z)=0$ where
$F(z)=z-\frac{(1-q)^{2}(1-z)}{1-q^{2}/b-z}-\frac{(b-1)q^{2}}{b^{2}(1-z)}.$ (24)
If $\rho_{R}=1/(b+1)$ then all edges are equivalent in a $b+1$ regular tree and the roots $z^{*}$ are $1/(b+1),b/(b+1)$. If $q=1$ (no useable red edges) then the problem reverts to a $b$-regular tree, in which case the roots $z^{*}$ of $F(z)$ are $1/b,(b-1)/b$.
If $b=2,q=1$, $z=1/2$ is the only root to $F(z)=0$. Ignoring this case which corresponds to a random walk on the line, we will assume that $q<1$ and $b\geq 2$. Let $z_{0}=1-q^{2}/b$. In the interval $I_{0}=[0,z_{0}]$, $F(0)<0$, $F(z_{0}^{-})=-\infty$. It can be checked that $F(z)$ is concave on $I_{0}$. In the interval $I_{1}=[z_{0},1]$, $F(z_{0}^{+})=\infty$, $F(1)=-\infty$ and $F(z)$ is monotone decreasing. So there are between one and three roots in $[0,1]$ depending on what happens in $I_{0}$. We will establish that if $b\geq 2,\,q<1$, then $F(1/2)>0$ which implies there are two roots in $I_{0}$. We have
@@
It can be checked that $\partial F/\partial b>0$ for any fixed $q$, so as $F(1/2)$ is monotone increasing with $b$, and $b=2$ has the largest range of $q$, (i.e., $12/3\leq q<$ 1), we check this case. After some simplification we obtain that if $b=2$,
@@
We solve $F(z)=0$, where $F(z)$ is given by (24), in the simplest case; namely $r=1$, and $b=2$. Thus $\rho_{R}+2\rho_{B}=1$. Put $q=1-\rho_{R}=2\rho_{B}$, and $w=1-z$ to give, after some rearrangement, $F(w)=N(w)/D(w)$. Here the numerator $N(w)$ is
@@
$\theta=\frac{1}{2}\left(\frac{1}{w+q/2}+\frac{1}{w-q/2}\right),$

### v3

@@ 1 Introduction
The cover time of a connected graph is the maximum over the start vertex of the expected time for a simple random walk to visit every vertex of the graph. There is a large literature on this subject see for example [1], [13], [14], and including [4][8] which give precise estimates of cover time for various models of random graphs.
@@
(Smooth walks) Let $G$ be the union of a random $r$-regular graph with red edges and a random $b$-regular graph with blue edges where $r\geq 1$ and $b\geq 2$.
@@
    If $b\geq 3$ and $\alpha<1$ constant, then w.h.p. the cover time of an $\alpha$-constrained smooth random walk on $G$ with budget $\gamma(\alpha)n\log n$ where $\gamma(\alpha)=\alpha\sigma_{RB}r/(r+b)$ is asymptotically equal to $C_{G}(R,B)=(\alpha\sigma_{RB}+(1-\alpha)\sigma_{B})n\log n$.
@@
In a flip walk there is no hard bound on the number of times we use a red edge, only a stochastic one. At any step, let $\rho_{R}$ be the probability of transition down a given red edge adjacent to the current vertex and $\rho_{B}$ the probability of transition down a given blue edge. Thus the probability the next transition is red is $r\rho_{R}$ etc. Examples include the following. Let $\rho_{R}=\alpha/r$, and $\rho_{B}=(1-\alpha)/b$ in which case the probability of transition along a red edge at any step is $\alpha$, and a blue edge is $1-\alpha$.
@@
    Suppose $b\geq 3$, and the budget is $(1-\varepsilon)\sigma_{RB}n\log n$. The cover time of an oblivious walk on $G$ is asymptotically equal to $(\alpha\sigma_{RB}+(1-\alpha)\sigma_{B})n\log n$ w.h.p., as given in (1) where $\alpha=\tfrac{1-\varepsilon}{r(r+b)}$.
@@
    Suppose that $b=2,r\geq 1$, and the blue edges span a Hamilton cycle. If the budget is at most $(1-\varepsilon)(r+1)/rn\log n$, the cover time of an oblivious walk on $G$ is asymptotic to $C_{G}(R,B)\sim n^{2}/2$, w.h.p.
@@
The main tool in our proofs will be the First Time Visit Lemma, a technique which will allow us to make very precise estimates of cover time in certain classes of graphs. The technique was introduced in [4], and subsequently refined in [7], and in [8] where various technical conditions were removed. In outline, the approach is to bound the cover time above and below using (7) of Lemma 4. Much of the work is in constructing the lower bound on the cover time. On regular graphs the walks we consider have uniform stationary distribution, and are rapidly mixing. The unvisit probability (7) has a parameter $p_{v}\sim\pi_{v}/R_{v}$ (see (6)) where $R_{v}$ is the expected number of returns to vertex $v$ during the mixing time. On regular graphs a simple random walk has a uniform stationary distribution, and is rapidly mixing on random $r$-regular graphs provided the vertex degree $r$ is at least three.
@@ 2 Background material
We use $A_{n}\sim B_{n}$ to denote $A_{n}=(1+o(1))B_{n}$ and thus $\lim_{n\rightarrow\infty}A_{n}/B_{n}=1$. We use $\omega$ to denote a quantity which tends to infinity with $n$ more slowly than any other variables in the given expression. The expression $f(n)\ll g(n)$ indicates $f(n)=o(g(n))$. The inequality $A\lesssim B$ which stands for $A\leq(1+o(1))B$ is used to unclutter notation in some places. A sequence of events ${\mathcal{E}}_{n}$ occurs with high probability, (w.h.p.), if $\lim_{n\to\infty}\mathbb{P}({\mathcal{E}}_{n})=1$.
Chernoff Bounds
We use the following versions of the Chernoff bounds, where $Bin(n,p)$ denotes a binomial random variable.
$\mathbb{P}(Bin(n,p)\geq(1+\varepsilon)np)$ $\leq e^{-\varepsilon^{2}np/3}\qquad\text{for $0\leq\varepsilon\leq 1$}.$ (2)
$\mathbb{P}(Bin(n,p)\leq(1-\varepsilon)np)$ $\leq e^{-\varepsilon^{2}np/2}\qquad\text{for $0\leq\varepsilon\leq 1$}.$ (3)
$\mathbb{P}(Bin(n,p)\geq\alpha np)$ $\leq\left(\frac{e}{\alpha}\right)^{\alpha np}\qquad\text{for $\alpha>0$}.$ (4)
@@ 2.1 First Time Visit Lemma
We denote the steady state probability of a simple random walk on $G=(V,E)$ by $\pi_{G}(v)$ for $v\in V$. We let $\mathbb{P}_{u,G}^{(t)}(v)$ denote the probability that a simple random walk on $G$ started at $u$ is at $v$ after $t$ steps. We then define a mixing time $T_{G}$ by
@@
For $t\geq 0$, let ${\mathcal{A}}_{t}(v)$ be the event that ${\mathcal{W}}_{u}$ does not visit $v$ in steps $T_{G},T_{G}+1,\ldots,t$. The vertex $u$ will have to be implicit in this definition. We will use the following version, it can be found in [7]:
@@
Let $R_{v}\geq 1$ is the expected number of visits by ${\mathcal{W}}_{v}$ to $v$ in the time interval $0,T_{G}]$. Suppose that $T_{G}\pi_{v}=o(1)$ and $R_{v}=O(1)$. Let
@@ 3.1 Proof of Theorem 1(a)
In this section $G$ is the union of a random $r$-regular graph with red edges, and a random $b$-regular graph with blue edges, where $r\geq 1$ and $b\geq 3$. We prove that if the budget is $\gamma(\alpha)n\log n$, where $\gamma(\alpha)=\alpha\sigma_{RB}r/(r+b)$, then w.h.p. the cover time of an $\alpha$-constrained smooth random walk on $G$ is asymptotically equal to $C_{G}=(\alpha\sigma_{RB}+(1-\alpha)\sigma_{B})\,n\log n$.
It is shown in [4] that w.h.p. the unconstrained cover time of $G$ is $\widehat{C}_{G}\sim\sigma_{RB}\,n\log n$. Thus without any restriction on the choice of the next edge, a walk of length $\widehat{C}_{G}$ would cross a red edge $\sim\frac{r}{r+b}\sigma_{RB}\,n\log n$ times. It follows that, for $\alpha<1$, w.h.p. we will not be able to complete the walk within the budget without having to avoid red edges a significant number of times.
If the phases are less than a mixing time in length ($o(T_{G})$), so that the walk is not smooth, there are some technical difficulties that we have not yet managed to overcome.
@@
There does not seem to be a compelling reason for us to only consider smooth walks. If we allow the walk to switch arbitrarily between allowing red edges and not allowing them, then because $b\geq 3$ there are no problems with the walk mixing. This is because the blue subgraph is an expander and the steady state probabilities are $1/n$ regardless. The problem arises because the $R_{v}$ depend heavily on how we switch between the two graphs.
@@ 3.1.1 Upper bound
Suppose now that we start a walk at vertex $u$ and that we alternate between red/blue phases and blue phases. Take $t_{0}=0$, and suppose we change phase at times $t_{1},t_{2},\ldots,t_{i},\ldots,t_{m}$ where $m=o(n^{2})$. Assume that whenever $i$ is even we start a red/blue phase at $t_{i}$ and switch to a blue phase when $i$ is odd. At each step in a red/blue phase there is an $r/(r+b)$ chance of using a red edge. We also know that this walk will finish at least as early as one that only uses blue edges. This follows from Lemma 4 and the fact that s/(s1)s/(s-1) decreases as $s$ increases. We will assume that $t_{i}-t_{i-1}\geq\omega\log^{2}n$, where $\omega\log^{2}n\gg T_{G}$, for $1\leq i<m$. It follows that
$t_{RB}=\sum_{i\geq 0}(t_{2i+1}-t_{2i})\leq\left(1-\frac{1}{\log^{1/2}n}\right)\alpha\sigma_{RB}n\log n.$ (9)
Applying the Chernoff bound (2), we see that if $Z_{i}$ denotes the number of times a red edge is used in $t_{2i},t_{2i+1}]$, then where $\varepsilon=1/\log^{1/2}n$,
$\mathbb{P}(Z_{R}\geq\gamma(\alpha)n\log n)\leq\sum_{i=0}^{m}\mathbb{P}\left(Z_{i}\geq\frac{(1+\varepsilon)(t_{2i+1}-t_{2i})r}{r+b}\right)\leq\sum_{i=0}^{m}\exp\left\{-\frac{\varepsilon^{2}r\omega\log^{2}n}{r+b}\right\}=o(1).$
@@ 3.1.2 Lower bound
For the lower bound we let $t_{1}=\tfrac{(1-\varepsilon)\alpha(r+b)}{(b-2)(r+b-1)}n\log n$ and argue that $C_{u}\geq t_{1}$ w.h.p. Proceeding as in [4], for any vertex $u$, we can find a set of vertices $S$ such that at time $t_{1}$, the probability the set $S$ is covered by the walk ${\mathcal{W}}_{u}$ tends to zero. Hence $T_{G}(u)>t_{1}$ w.h.p. which together with (16) completes the proof of Theorem 1(a).
We construct $S$ as follows. Let $\sigma$ be given by (8). Let $S\subseteq[n]$ be some maximal set of locally tree-like vertices all of which are at least distance $2\sigma+1$ apart. Thus $|S|\geq(n-r^{3\sigma})/r^{(2\sigma+1)}$.
@@
$\mathbb{P}(S(t_{1})\neq 0)\geq\frac{(\mathbb{E}|S(t_{1})|)^{2}}{\mathbb{E}(|S(t_{1})|^{2})}=\frac{1}{\frac{{\bf E}|S_{t_{1}}|(|S_{t_{1}}|-1)}{({\bf E}|S(t_{1})|)^{2}}+(\mathbb{E}|S_{t_{1}}|)^{-1}}=1-o(1).$
@@
There is a natural measure-preserving mapping from the set of walks in $G$ which start at $u$ and do not visit $v$ or $w$, to the corresponding set of walks in $\Gamma$ which do not visit $Z$. Thus the probability that ${\mathcal{W}}_{u}$ does not visit $v$ or $w$ in the first $t$ steps is equal to the probability that a random walk $\widehat{{\mathcal{W}}}_{u}$ in $\Gamma$ which also starts at $u$ does not visit $Z$ in the first $t$ steps.
@@ 4 Proof of Theorem 2
where $E(S:S^{c})$ is the number of edges between $S$ and $S^{c}$ and ${\Phi}_{SRW}$ refers to the conductance of a simple random walk.
As $G$ satisfies P5, the conductance ${\Phi}^{*}_{SRW}$ of a simple random walk on $G$ is constant for some $0<c<1$. Assuming $\min(\rho_{R},\rho_{B})$ is also a positive constant, the conductance ${\Phi}^{*}_{PB}$ of the flip walk is constant, i.e.,
@@
Consider an infinite $d$-regular tree $T_{d}$, where $r+b=d$ and in which each vertex has $r$ red and $b$ blue edges. For discussion we regard $T_{d}$ as rooted at a designated vertex $v$, with all edges nominally oriented outward from $v$. With this orientation, a transition away from $v$ (resp. $u$) at a vertex $u$, is a transition $uw$ down an out-edge of $u$. Relative to this orientation, a vertex other than $v$ is red if its in-edge is red, and blue if its in-edge is blue. With the exception of $v$ this induces an alternating pattern of red and blue vertices in $T_{d}$, in which e.g., a red vertex has (r1)(r-1) red out-edges, and $b$ blue out-edges.
@@
[…]
and define $\phi_{RB},\phi_{BB},\phi_{BR}$ similarly. Note that $\phi_{RR}=\phi_{BR}=\phi_{v,R}$. We denote this common value by $\phi_{R}$. For a red vertex $u$, let $\psi_{R}(u)$ be the probability the walk moves away from $v$ at $u$, and subsequently returns to $u$. Thus with $P(u,w)$ the colour dependent transition probability over the edge $uw$, $\phi(uw)$ the colour dependent return probability, and $N^{+}_{R}(u)$ the red out-neighbours of $u$ etc.,
@@
If $u$ is red, there are r1r-1 red out-neighbours, and $b$ blue out-neighbours, so
@@
As $r=1$ we have $\rho_{R}+b\rho_{B}=1$. Assume $b\geq 2$ and $\rho_{R}\leq\rho_{B}$ so that $0\leq\rho_{R}\leq 1/(b+1)$. Use (22) to get
$\frac{1}{1-\psi_{R}}=\frac{1-\psi_{B}}{1-\psi_{B}-b\rho_{B}^{2}},$
and substitute this in (23) giving
$\psi_{B}=\frac{\rho_{R}^{2}(1-\psi_{B})}{1-\psi_{B}-b\rho_{B}^{2}}+\frac{(b-1)\rho_{B}^{2}}{1-\psi_{B}}.$
Put $q=b\rho_{B}$ and $\rho_{R}=1-q$ where $b/(b+1)\leq q\leq 1$, and replace $\psi_{B}$ by $z$ in the above expression. We look for solutions to $F(z)=0$ where
$F(z)=z-\frac{(1-q)^{2}(1-z)}{1-z-q^{2}/b}-\frac{(b-1)q^{2}}{b^{2}(1-z)}.$ (27)
If $\rho_{R}=1/(b+1)$ then all edges are equivalent in a $b+1$ regular tree and the roots $z^{*}$ are $1/(b+1),b/(b+1)$. If $q=1$ (no useable red edges) then the problem reverts to a $b$-regular tree, in which case the roots $z^{*}$ of $F(z)$ are 1/b,(b1)/b1/b,(b-1)/b.
If $b=2,q=1$, $z=1/2$ is the only root to $F(z)=0$. Ignoring this case which corresponds to a random walk on the line, we will assume that $q<1$ and $b\geq 2$. Let $z_{0}=1-q^{2}/b$. In the interval $I_{0}=[0,z_{0}]$, $F(0)<0$, $F(z_{0}^{-})=$-$\infty$. It can be checked that $F(z)$ is concave on $I_{0}$. Indeed, $z$ and $\frac{(b-1)q^{2}}{b^{2}(1-z)}$ are concave on $[0,1)$ and we can write $\frac{1-z}{1-z-q^{2}/b}=-1-\frac{q^{2}/b}{1-z-q^{2}/b}$.
In the interval $I_{1}=[z_{0},1]$, $F(z_{0}^{+})=\infty$, $F(1)=$-$\infty$ and $F(z)$ is monotone decreasing. So there are between one and three roots in $[0,1]$ depending on what happens in $I_{0}$. We will establish that if $b\geq 2,\,q<1$, then $F(1/2)>0$ which implies there are two roots in $I_{0}$. We have
@@
It can be checked that $\partial F/\partial b>0$ for any fixed $q$, so as $F(1/2)$ is monotone increasing with $b$, and $b=2$ has the largest range of $q$, (i.e., $2/3\leq q<1$), we check this case. After some simplification we obtain that if $b=2$,
@@
We solve $F(z)=0$, where $F(z)$ is given by (27), in the simplest case; namely $r=1$, and $b=2$. Thus $\rho_{R}+2\rho_{B}=1$. Put $q=1-\rho_{R}=2\rho_{B}$, and w=1zw=1-z to give, after some rearrangement, $F(w)=N(w)/D(w)$. Here the numerator $N(w)$ is
@@
$\theta=\frac{1}{2}\left(\frac{1}{w+q/2}+\frac{1}{w-q/2}\right).$